OpenStudy (twizttiez):

Which functions have a positive rate of change and which have a negative rate of change? Column A Column B 1. 2x + 3y = −9 2. y = 3x − 6 3. x y −1 9 0 7 1 5 4. x y 2 −8 3 −7 4 −6

OpenStudy (twizttiez):

Are these functions positive or negative?

OpenStudy (twizttiez):

@DanJS

OpenStudy (danjs):

rate of change is also known as , slope of the line , change in y coordinate vs a change in x coordinate

OpenStudy (danjs):

2x + 3y = −9 rearrange it , to y= m*x + b look at the sign on m, + or - rate of change

OpenStudy (twizttiez):

So negative?

OpenStudy (danjs):

right

OpenStudy (danjs):

The second one is already in form y = m*x+b look at the sign on m

OpenStudy (twizttiez):

Ok so 2 is positive?

OpenStudy (danjs):

yes, rate of change 3 y values per 1 x value

OpenStudy (twizttiez):

Ok now how do i solve the charts?

OpenStudy (danjs):

The next 2, they give you points on the line (x,y) and (x1,y1) calculate m = (y-y1) / (x-x1) that is the slope (rate of change )

OpenStudy (twizttiez):

Ok i think i get it

OpenStudy (danjs):

For 3) you can use 2 points (-1 , 9) (0 , 7) calculate difference in Y coordinates over difference in X coordinates

OpenStudy (twizttiez):

So it would be positive?

OpenStudy (twizttiez):

And 4 would be negative?

OpenStudy (danjs):

rate of change = (7 - 9) / (0 -(-1)) = -2 / 1 = -2

OpenStudy (twizttiez):

Oh so opposite?

OpenStudy (twizttiez):

3 is positive and 4 is negative?

OpenStudy (twizttiez):

I mean negative

OpenStudy (twizttiez):

And 4 is positive lol sorry

OpenStudy (danjs):

For 3) 2 points (-1 , 9) and (0,7) \[m = \frac{ y - y1 }{ x - x1 } = \frac{ 7 - 9 }{ 0 - (-1) } = \frac{ -2 }{ 1 } = -2\] negative rate of change

OpenStudy (danjs):

do the same thing for 4

OpenStudy (twizttiez):

Ok

OpenStudy (twizttiez):

I got a positive 1 for 4 is that right??

OpenStudy (twizttiez):